alones (alones):

Solve .....this ......quadratic equation -----------------------------------------------------------------------------------(using the quadratic formula.)

OpenStudy (usukidoll):

@johnweldon1993 here it is

alones (alones):

Solve this quadratic equation using the quadratic formula. 2x^2 - 2x = 1 check

OpenStudy (johnweldon1993):

XD you did not! @AloneS XD Well I'm going to sit quietly in the background while you list the steps for me >:D

alones (alones):

O.M.G. NOOOOO

alones (alones):

Damn i'm dead

OpenStudy (usukidoll):

it's either that or prove that f+g has a limit at x_0

alones (alones):

usiki o.m.g why did u do itttttt >.>

alones (alones):

Okay a,b,c,d a=2x^2 b=2x c=1?

alones (alones):

let me look , i forgo tthe steps again.brb

OpenStudy (johnweldon1993):

I'm going to post them 1 last time :) You have an equation in the form \(\large ax^2 + bx + c = 0\) 1) Find your coefficients. a = ? b = ? c = ? 2) Write down the quadratic formula \(\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) 3) Plug in your values 4) simplify if possible

alones (alones):

OOooooh thank you *bows*

alones (alones):

Yo don't delete it tho

OpenStudy (usukidoll):

ok * secretly erases*

OpenStudy (johnweldon1993):

I won't as long as you get it 100% correct >:D

OpenStudy (johnweldon1993):

1 wrong move...POOF...gone >:D * we are in the fun part of openstudy now lmao

Nnesha (nnesha):

i think the first step should be rewrite the equation in ax^2+bx+C=0 quadratic form 2) Find your coefficients. a = ? b = ? c = ? 3) Write down the quadratic formula \(\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) 4) Plug in your values 5) simplify if possible

alones (alones):

whooa usiki dont' let me look like a slave there. LOL WELDOONE where is it!! I WANNAA SEE IT Thansk nnesha well done can earase it cx

Nnesha (nnesha):

?

alones (alones):

\[\frac{ -2\sqrt{(-2)^2-4(2)(-1)} }{ 2(2) }\]

alones (alones):

okay what's next *scrathes head*

alones (alones):

Okay okay so \[\frac{ 2+\sqrt{12} }{ 4 }\]

Nnesha (nnesha):

lets play game compare this \[\frac{ -2\sqrt{(-2)^2-4(2)(-1)} }{ 2(2) }\] with \(\large \rm \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) can you find the difference ?

alones (alones):

yay! difference in what?o.o

Nnesha (nnesha):

nvm i think you don't know how to type plus/minus sign `/pm`

alones (alones):

yea yeaa i do liek this

alones (alones):

Oh damn it didn't post\[nesha \pm human\]

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @AloneS \[\frac{ -2 \sqrt{(-2)^2-4(2)(-1)} }{ 2(2) }\] \(\color{blue}{\text{End of Quote}}\) \[\frac{ \color{red}{-}2 \pm \sqrt{(-2)^2-4(2)(-1)} }{ 2(2) }\] \(\color{blue}{\text{Originally Posted by}}\) @AloneS Okay okay so \[\frac{ 2+\sqrt{12} }{ 4 }\] \(\color{blue}{\text{End of Quote}}\) what happened to that negative sign ?

Nnesha (nnesha):

what's `b` value in that equation ?? :)

alones (alones):

It turned in to postitive :D lol b=-2 and i got my answer \[x=\frac{ 1\pm \sqrt{32} }{ 24 }\]

Nnesha (nnesha):

`It turned in to postitive :D` right so it should be \[\frac{ \color{red}{-}(\color{red}{-2}) \pm \sqrt{(\color{Red}{-2})^2-4(2)(-1)} }{ 2(2) }\] b is negative and there is another negative sign in quadratic formula

Nnesha (nnesha):

you're tired..right?

alones (alones):

Oh ye.. I am tbh.

alones (alones):

waaait show it again

Nnesha (nnesha):

i think you mean tto type 4 at the denominator ? \[x=\frac{ 1\pm \sqrt{32} }{ 4 }\]

alones (alones):

Oh lol yeaa i didn't reduce it thansk<3

Nnesha (nnesha):

\[\frac{ \color{red}{-}(\color{red}{-2}) \pm \sqrt{(\color{Red}{-2})^2-4(2)(-1)} }{ 2(2) }\] simplify this ..again

alones (alones):

Oh damn again? ;~;

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) AloneS It turned in to postitive :D lol b=-2 and i got my answer \[x=\frac{ 1\pm \sqrt{32} }{ 24 }\] \(\color{blue}{\text{End of Quote}}\) yea or can you please explain what happened to 2 at the numerator ? how did you get 32 ?

alones (alones):

Nooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooo*rips the hair out*

alones (alones):

Oh fine i'll explain i mulitplied (-2)by square root 2, than i mulitplied 4*1 which gave me 4, and than multiplied 4*4 which is 16 and i ended up 32

Nnesha (nnesha):